OpenStudy (anonymous):

what is the y-coordinate of the point of intersection for the two lines below? -6x+7y=20 2x-3=4 help pleaseee?!

OpenStudy (anonymous):

I think your second equation is missing a y.

OpenStudy (anonymous):

yes it is sorry its 3y

OpenStudy (anonymous):

Ok, so if you want to find the y coordinate you have to eliminate the x's. What can you multiply one of your equations by to make the coefficient of the x variable have the same magnitude but opposite sign?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Multiply the second equation by 3.

OpenStudy (anonymous):

Then add the equations.

OpenStudy (anonymous):

so then itll x-2=8 would i add two to each side then devide by x?

OpenStudy (anonymous):

Huh? You shouldn't have any x's

OpenStudy (anonymous):

Show me the steps you took.

OpenStudy (anonymous):

-6x+7y=20 3(2x-3y=4) ________________ -6x+7y=20 6x-9y=12 x-2=8 2 2 x=10?

OpenStudy (anonymous):

6x - 6x = x?

OpenStudy (anonymous):

I'm pretty sure -6x + 6x = 0. And I'm pretty sure 7y - 9y = -2y

OpenStudy (anonymous):

So try it again?

OpenStudy (anonymous):

y=-4?

OpenStudy (anonymous):

Very close. You have the sign right

OpenStudy (anonymous):

Show your steps again?

OpenStudy (anonymous):

my practice test packet says the answer is -16 i can't see where its coming from :/ im confused

OpenStudy (anonymous):

You are really close!

OpenStudy (anonymous):

-6x+7y=20 + 3(2x-3y=4) ________________ -6x+7y=20 + 6x -9y=12 ========== This was all correct

OpenStudy (anonymous):

Just add them up better.

OpenStudy (anonymous):

i just realized what i did i subtracted 12 from 20 instead of adding

OpenStudy (anonymous):

There ya go.

OpenStudy (anonymous):

thank you so much! so do i do that everytime im trying to find the y-coordinate?

OpenStudy (anonymous):

To find the y you eliminate the x.

OpenStudy (anonymous):

You could also use substitution.

OpenStudy (anonymous):

But it amounts to the same thing.

OpenStudy (anonymous):

thanks!